# Multiscale scattered data analysis in samplet coordinates

### v1

@@ Abstract.
We study multiscale scattered data interpolation schemes for globally supported radial basis functions, with a focus on the Matérn class. The multiscale approximation is constructed through a sequence of residual corrections, where radial basis functions with different lengthscale parameters are employed to capture varying levels of detail. To apply this approach to large data sets, we suggest to represent the resulting generalized Vandermonde matrices in samplet coordinates. Samplets are localized, discrete signed measures exhibiting vanishing moments and allow for the sparse approximation of generalized Vandermonde matrices issuing from a vast class of radial basis functions. Given a quasi-uniform set of $N$ data sites, and local approximation spaces with geometrically decreasing dimension, the full multiscale system can be assembled with cost $\mathcal{O}(N\log N)$. We prove that the condition numbers of the linear systems at each level remain bounded independent of the particular level, allowing us to use an iterative solver with a bounded number of iterations for the numerical solution. Hence, the overall cost of the proposed approach is $\mathcal{O}(N\log N)$. The theoretical findings are accompanied by extensive numerical studies in two and three spatial dimensions.
@@ 1. Introduction
Interpolation using radial basis functions (RBFs) is a widely recognized technique for fitting functions based on scattered data in $\mathbb{R}^{d}$, see [10, 13, 14] for some surveys. RBF interpolation allows for the construction of approximation spaces in arbitrary dimensions and with varying levels of smoothness. These spaces are well known to provide high-quality approximations, see [11, 25]. However, the actual computation of an RBF interpolant usually entails a significant computational cost and the corresponding linear system is typically ill-conditioned. Even worse, as the convergence order increases, the associated problems become successively ill-conditioned. Moreover, using globally supported RBFs in interpolation leads to densely populated matrices, making the problem computationally infeasible for larger datasets. To address the latter, techniques such as interpolation using compactly supported RBFs have been developed [19, 24]. Unfortunately, even interpolation with these kernels becomes either computationally too expensive, if unscaled kernels are used, or do not even lead to convergence, if the support is scaled proportional to the fill-distance. Motivated by solving this trade-off principle, the works [3, 5, 9, 21], among many others, leverage the idea of a multiscale method where approximations on different scales are computed in an residual correction scheme. To date, the most general convergence theorems for multiscale interpolation using positive definite, compactly supported RBFs have been developed in [21]. The advantage of the multiscale method is that the condition numbers of the resulting linear systems at each level remain independent of the particular level. Thus achieving linear convergence in the number of used levels with relatively low cost.
@@
In this article, we study the RBF interpolation problem using globally supported RBFs. We address the challenge of dealing with densely-populated interpolation matrices by employing samplet compression, a technique introduced in [7]. Samplets are discrete signed measures constructed such that all polynomials up to a certain degree vanish. This vanishing moment property allows for the compression of generalized Vandermonde matrices issuing from asymptotically smooth RBFs, i.e., they behave like smooth functions apart from zero. Hence, In samplet coordinates, the resulting interpolation matrices become quasi-sparse, which means that they can be compressed such that only a sparse matrix remains. Furthermore, samplets allow for a respective sparse matrix algebra, see [8], and provide a meaningful interpretation of sparsity constraints for scattered data, see [2]. However, as the samplet transform of a given generalized Vandermonde matrix is an isometry up to the compression error, it does not alleviate the ill-conditioning. Therefore, to mitigate issues related to ill-conditioning, we adopt a multiscale interpolation approach that exploits scaled versions of the same RBF, which, in our case, remains globally supported. Hence, recovering the results from [21] for the class of Matérn kernels.
@@
The remainder of this article is organized as follows. In Section 2, we revisit the classical multiscale interpolation algorithm, detailing the relevant subspaces and the class of radial basis functions considered in this work. In Section 3 we carry over a known result for compactly supported functions to globally supported RBFs: The condition number of the kernel matrix can be bounded uniformly in the number of sites, if we apply the appropriate scaling. Section 4 provides a brief introduction of samplets, with a focus on the key properties of the basis transformation and compression. In Section 5, we present the multiscale interpolation algorithm within the samplet basis and show an example of the resulting generalized Vandermonde matrix pattern. Finally, in Section 6, we provide numerical tests in both two and three spatial dimensions. Concluding remarks are stated in Section 7.
@@ 2. Classical multiscale method
Let $\Omega\subseteq\mathbb{R}^{d}$ be a domain and $X=\{{\boldsymbol{x}}_{1},\ldots,{\boldsymbol{x}}_{N}\}\subset\Omega$ be a set of data sites with cardinality $N=\#X$. Moreover, let $f\in C(\Omega)$ be the target function generating the data $f_{1},\dots,f_{N}$. Associated to the set of sites $X$ there are two measures, the fill-distance of $X$ in $\Omega$ and the separation radius of $X$,
$h_{X,\Omega}$ $\mathrel{\mathrel{\mathop{:}}=}\sup_{{\boldsymbol{x}}\in\Omega}\min_{{\boldsymbol{x}}\in X}\|{\boldsymbol{x}}-{\boldsymbol{x}}_{j}\|_{2}$
$q_{X}$ $\mathrel{\mathrel{\mathop{:}}=}\frac{1}{2}\min_{j\neq k}\|{\boldsymbol{x}}_{j}-{\boldsymbol{x}}_{k}\|_{2}.$
@@
We call $X$ quasi-uniform if there is a constant $c_{\text{qu}}>0$ such that
$q_{X}\leq h_{X,\Omega}\leq c_{\text{qu}}q_{X}.$
Our goal is to interpolate or approximate the target function by means of radial basis functions (RBF) using only the given data $({\boldsymbol{x}}_{1},f_{1}),\ldots,({\boldsymbol{x}}_{N},f_{N})$. A function $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ is said to be radial, if and only if there exists a univariate function $\phi\colon[0,\infty)\to\mathbb{R}$ such that $\Phi({\boldsymbol{x}})=\phi(\|{\boldsymbol{x}}\|_{2})$. The radial function $\Phi$ is strictly positive definite, if and only if there holds for any choice of distinct points ${\boldsymbol{\xi}}_{1},\ldots,{\boldsymbol{\xi}}_{n}$, $n\in\mathbb{N}$, that
${\boldsymbol{\alpha}}^{\intercal}[\Phi({\boldsymbol{\xi}}_{i}-{\boldsymbol{\xi}}_{j})]_{i,j=1}^{n}{\boldsymbol{\alpha}}>0\quad\text{for all }{\boldsymbol{\alpha}}\in\mathbb{R}^{n}\setminus\{{\boldsymbol{0}}\}.$
In this case, the kernel
$K({\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}})\mathrel{\mathrel{\mathop{:}}=}\Phi({\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j})$
@@
is the reproducing kernel of a uniquely determined reproducing kernel Hilbert space, its native space, that we denote by $\big{(}\mathcal{N}_{\Phi},\langle\cdot,\cdot\rangle_{\mathcal{N}_{\Phi}}\big{)}$. In this context, we will use the function $\Phi$ and its associated kernel $K$ synonymously. A popular class of strictly positive RBFs are the Matérn kernels or Sobolev splines $\Phi_{\nu}\colon\mathbb{R}^{d}\to\mathbb{R}$, dependent on a hyper-parameter $\nu>d/2$. These kernels are defined via the radial function
(1) $\phi_{\nu}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}r^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}(r),\quad r\geq 0,$
@@
where $\Gamma$ is the Riemann Gamma function and $K_{\beta}$ is the modified Bessel function of the second kind, see [12] for example.
It is well-known that the Fourier transform of this radial function behaves like
(2) $\widehat{\Phi}(\boldsymbol{\omega})=(1+\|\boldsymbol{\omega}\|_{2}^{2})^{-\nu-\frac{d}{2}},\quad\boldsymbol{\omega}\in\mathbb{R}^{d},$
@@
and that, in turn, the native space of the corresponding Matérn kernel is norm-equivalent to the classical Sobolev space $H^{\nu+\frac{d}{2}}(\mathbb{R}^{d})$, see [20, Theorem 6.13] for example.
In this instationary setting it is known that the interpolant
$s\in\mathcal{S}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X\}\subset\mathcal{N}_{\Phi}$
@@
converges towards the target function $f$ as $h_{X,\Omega}\to 0$. However, this improved approximation comes with higher and higher numerical cost for the iterative solution due to the ill-conditioning. To alleviate this cost, we can rescale the kernel, i.e., we introduce the parameter $\delta>0$ and define
@@
In the stationary setting we couple $\delta$ proportional to $h_{X,\Omega}$, i.e., there is a coupling parameter $\eta>0$ such that $\delta=\eta h_{X,\Omega}$. In this case we can control the numerical costs to compute the interpolant but cannot expect convergence. This is known as the trade-off principle.
@@
where we refer to the subscript of $X_{\ell}$ as level with cardinalities $N_{\ell}$ and fill-distances $h_{\ell}\mathrel{\mathrel{\mathop{:}}=}h_{X_{\ell},\Omega}$. Although in many applications the sets will be nested, for the multiscale method itself, this is not necessary. We assume that the cardinalities of the sets $X_{\ell}$ are increasing with the level, i.e., $N_{\ell}\leq N_{\ell+1}$ or, in other words, $h_{\ell}\geq h_{\ell+1}$. We quantify this relation in Theorem 2.1. Now, we use the stationary setting on each level $\ell$. We have level-dependent $\delta_{\ell}>0$ and level-dependently rescaled kernel $K_{\ell}$ given by
$K_{\ell}(\boldsymbol{x},\boldsymbol{y})\mathrel{\mathrel{\mathop{:}}=}\Phi_{\ell}(\boldsymbol{x}-\boldsymbol{y})=\delta_{\ell}^{-d}\Phi\left(\frac{\boldsymbol{x}-\boldsymbol{y}}{\delta_{\ell}}\right),\quad\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$
@@
also called local approximation space.
With the multiscale sequence of points and the associated kernels, we can decompose the original approximation problem into $L$ steps using a telescoping sum
@@
The intuition is that at the $\ell$-th step we compute the approximation to the error function or residual of level $\ell-1$, i.e.,
(6) $f-(w_{1}+\cdots+w_{l-1})$
in the space $\mathcal{S}_{\ell}$, given as
(7) $w_{\ell}({\boldsymbol{x}})=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}({\boldsymbol{x}}-{\boldsymbol{x}}_{i}),\quad\boldsymbol{x}\in\Omega,$
with a coefficient vector ${\boldsymbol{c}}_{\ell}\in\mathbb{R}^{N_{\ell}}$.
From this, we obtain the multiscale decomposition
$s_{L}\mathrel{\mathrel{\mathop{:}}=}(w_{1}+\cdots+w_{L})|_{X}=f|_{X},$
which means that $s_{L}\in\cup_{\ell=1}^{L}\mathcal{S}_{\ell}$ matches $f$ at each of the data sites in $X$.
In practice, the multiscale approximation process begins with a coarse set of points, where the approximation problem (7) is solved. Next, the error function (6) is computed, and finally, the latter is approximated at a finer scale. Interpolating the residual at level $\ell$ can be written in the following form:
$\boldsymbol{K}_{\ell,\ell}\boldsymbol{c}_{\ell}=\boldsymbol{f}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}_{\ell,\ell^{\prime}}\boldsymbol{c}_{\ell^{\prime}},$
where
$\boldsymbol{K}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\big{[}\Phi_{\ell^{\prime}}\big{(}{\boldsymbol{x}}_{i}^{(\ell)}-\boldsymbol{x}_{j}^{(\ell^{\prime})}\big{)}\big{]}_{i,j},\quad\ell,\ell^{\prime}=1,\ldots,L.$
Therefore, the algorithm to solve the multiscale interpolation consists of solving the following triangular system, whose entries are matrices of different sizes:
(8) $\left[\begin{array}[]{ccccc}\boldsymbol{K}_{1,1}&&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&&\\ \boldsymbol{K}_{3,1}&\boldsymbol{K}_{3,2}&\boldsymbol{K}_{3,3}&&\\ \vdots&\vdots&\cdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L-1}&\boldsymbol{K}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \boldsymbol{c}_{3}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \boldsymbol{f}_{3}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]$
The resulting matrix is in general a densely-populated populated lower triangular matrix as a result of the choice of globally supported RBFs for the approximation at each level. The samplet compression plays a crucial role here, rendering the matrix quasi-sparse. This technique is introduced in Section 4, and the linear system is reformulated in the samplet basis in Section 5.
There holds the following convergence result, see, e.g., [21, 22].
@@
for $\ell=1,2,\dots$ with fixed $\mu\in(0,1)$, $c\in(0,1]$ and $h_{1}$ sufficiently small. Let $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel for $H^{s}(\mathbb{R}^{d})$, i.e., its Fourier transform satisfies
$c_{1}(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-s}\leq\widehat{\Phi}({\boldsymbol{\omega}})\leq c_{2}(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-s},\quad{\boldsymbol{\omega}}\in\mathbb{R}^{d},$
and let $\Phi_{\ell}$ be the rescaled kernel as in (3) with scale factor $\delta_{\ell}=\eta h_{\ell}$, $\eta>0$. Let $f\in H^{s}(\Omega)$. Then there exist constants $C$ and $C_{1}>0$ such that
$\|f-s_{L}\|_{L_{2}(\Omega)}\leq C(C_{1}\mu^{s}+C_{1}\eta^{-s})^{L}\|f\|_{H^{s}(\Omega)}.$

### v2

@@ Abstract.
We study multiscale scattered data interpolation schemes for globally supported radial basis functions with focus on the Matérn class. The multiscale approximation is constructed through a sequence of residual corrections, where radial basis functions with different lengthscale parameters are combined to capture varying levels of detail. We prove that the condition numbers of the the diagonal blocks of the corresponding multiscale system remain bounded independently of the particular level, allowing us to use an iterative solver with a bounded number of iterations for the numerical solution. Employing an appropriate diagonal scaling, the multiscale system becomes well conditioned. We exploit this fact to derive a general error estimate bounding the consistency error issuing from a numerical approximation of the multiscale system. To apply the multiscale approach to large data sets, we suggest to represent each level of the multiscale system in samplet coordinates. Samplets are localized, discrete signed measures exhibiting vanishing moments and allow for the sparse approximation of generalized Vandermonde matrices issuing from a vast class of radial basis functions. Given a quasi-uniform set of $N$ data sites, and local approximation spaces with exponentially decreasing dimension, the samplet compressed multiscale system can be assembled with cost $\mathcal{O}(N\log^{2}N)$. The overall cost of the proposed approach is $\mathcal{O}(N\log^{2}N)$. The theoretical findings are accompanied by extensive numerical studies in two and three spatial dimensions.
@@ 1. Introduction
Interpolation using radial basis functions (RBFs) is a widely recognized technique for fitting functions based on scattered data in $\mathbb{R}^{d}$, see [17, 20, 21] for some surveys. RBF interpolation allows for the construction of approximation spaces in arbitrary dimensions and with varying levels of smoothness. These spaces are well known to provide high-quality approximations, see [18, 37]. Even so, the computation of an RBF interpolant usually entails a significant computational cost and the corresponding linear system is typically ill-conditioned. Even worse, as the convergence order increases, the associated problems become successively ill-conditioned. Moreover, using globally supported RBFs in interpolation leads to densely populated matrices, making the problem computationally infeasible for larger datasets. To address the latter, techniques such as interpolation using compactly supported RBFs have been developed [27, 36]. Unfortunately, even interpolation with these RBFs either becomes computationally too expensive, if unscaled kernels are used, or does not even lead to convergence, if the support is scaled proportional to the fill-distance. Motivated by balancing this trade-off principle, the works [4, 6, 15, 29], among many others, leverage the idea of a multiscale method where approximations on different scales are computed in a residual correction scheme. To date, the most general convergence theorems for multiscale interpolation using positive definite, compactly supported RBFs have been developed in [29]. The advantage of the multiscale method is that the condition numbers of the resulting linear systems at each level remain independent of the particular level.
An effective approach to address the computational challenges of large datasets is the greedy kernel method [33]. The greedy kernel method systematically selects a representative subset of the data to construct a sparse approximation. An alternative strategy involves deep kernel methods [3, 23, 34, 32]. These methods leverage deep neural networks to learn adaptive feature representations, thereby constructing kernels that are tuned to the intrinsic structure of medium to high-dimensional data. Further, kernel multigrid methods [35, 13, 7] have been developed for the efficient numerical solution of partial differential equations on complex domains such as manifolds or point cloud surfaces. These techniques use specialized kernels, such as polyharmonic or homogeneous kernels, along with localized Lagrange bases.
@@
In this article, we study the RBF interpolation problem using globally supported RBFs. We address the challenge of dealing with densely-populated interpolation matrices by employing samplet compression, a technique introduced in [9]. Samplets are discrete signed measures constructed such that all polynomials up to a certain degree vanish. This vanishing moment property allows for the compression of generalized Vandermonde matrices issuing from asymptotically smooth RBFs, i.e., they behave like smooth functions apart from zero. Hence, in samplet coordinates, the resulting interpolation matrices become quasi-sparse, which means that they can be compressed such that only a sparse matrix remains. Furthermore, samplets allow for a respective sparse matrix algebra, see [11], and provide a meaningful interpretation of sparsity constraints for scattered data, see [2]. Unlike approaches that rely on the specific kernel, samplets can be applied to a broad class of positive definite kernels.
However, as the samplet transform of a given generalized Vandermonde matrix is an isometry up to the compression error, it does not alleviate the ill-conditioning. Therefore, to mitigate issues related to ill-conditioning, we adopt a multiscale interpolation approach that exploits scaled versions of the same RBF, which, in our case, remains globally supported. Especially, we extend the results from [29] for compactly supported RBFs to the class of Matérn kernels. As a consequence, the diagonally scaled multiscale system becomes well conditioned. We exploit this fact to derive a general error estimate bounding the consistency error issuing from a numerical approximation of the multiscale system. The theoretical results are corroborated by extensive numerical studies in two and three spatial dimensions.
@@
The remainder of this article is organized as follows. In Section 2, we revisit the classical multiscale interpolation algorithm, detailing the relevant subspaces and the class of radial basis functions considered in this work. In Section 3 we carry over a known result for compactly supported functions to globally supported RBFs: The condition number of the generalized Vandermonde matrix can be bounded uniformly in the number of sites, if an appropriate scaling is applied. Section 4 derives a general approximation error estimate for the case that the block matrix in the multiscale system is numerically approximated. Section 5 provides a brief introduction of samplets, with a focus on the key properties of the basis transformation and compression. In Section 6, we present the multiscale interpolation algorithm in the samplet basis and provide error and cost estimates. Finally, in Section 7, we provide numerical tests in both two and three spatial dimensions. Concluding remarks are stated in Section 8.
Throughout this article, to avoid the excessive use of generic constants, we use the notation $A\sim B$ to indicate that there exists constants $\underline{c},\overline{c}>0$ independent of $A,B$, and any quantities they may depend on, such that $\underline{c}A\leq B\leq\overline{c}A$. In addition $C>0$ will always denote a generic constant, whose value can change depending on the context.
@@ 2. Classical multiscale method
Let $\Omega\subseteq\mathbb{R}^{d}$ be a domain and $X=\{{\boldsymbol{x}}_{1},\ldots,{\boldsymbol{x}}_{N}\}\subseteq\Omega$ be a set of data sites with cardinality $N\mathrel{\mathrel{\mathop{:}}=}\#X$. Moreover, let $f\in C(\Omega)$ be the function generating the data values $f_{1},\dots,f_{N}$ corresponding to $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{N}$. Associated to the set of data sites $X$, there are two measures, namely the fill-distance of $X$ in $\Omega$ and the separation radius of $X$, i.e.,
$h_{X,\Omega}\mathrel{\mathrel{\mathop{:}}=}\sup_{{\boldsymbol{x}}\in\Omega}\min_{{\boldsymbol{x}}\in X}\|{\boldsymbol{x}}-{\boldsymbol{x}}_{j}\|_{2},\quad q_{X}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{2}\min_{j\neq k}\|{\boldsymbol{x}}_{j}-{\boldsymbol{x}}_{k}\|_{2}.$
@@
We call $X$ quasi-uniform if there exists a constant $c_{\text{qu}}>0$ such that $q_{X}\leq h_{X,\Omega}\leq c_{\text{qu}}q_{X}.$
@@
Our goal is to interpolate or approximate the target function by means of RBFs only relying on the given data $({\boldsymbol{x}}_{1},f_{1}),\ldots,({\boldsymbol{x}}_{N},f_{N})$. A function $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ is said to be radial, if and only if there exists a univariate function $\phi\colon[0,\infty)\to\mathbb{R}$ such that $\Phi({\boldsymbol{x}})=\phi(\|{\boldsymbol{x}}\|_{2})$. The radial function $\Phi$ is strictly positive definite, if and only if there holds for any choice of distinct points ${\boldsymbol{\xi}}_{1},\ldots,{\boldsymbol{\xi}}_{n}$, $n\in\mathbb{N}$, that ${\boldsymbol{\alpha}}^{\intercal}[\Phi({\boldsymbol{\xi}}_{i}-{\boldsymbol{\xi}}_{j})]_{i,j}{\boldsymbol{\alpha}}>0\ \text{for all }{\boldsymbol{\alpha}}\in\mathbb{R}^{n}\setminus\{{\boldsymbol{0}}\}.$ In this case, the kernel $K({\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}})\mathrel{\mathrel{\mathop{:}}=}\Phi({\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j})$ is the reproducing kernel of a uniquely determined reproducing kernel Hilbert space, its native space, that we denote by $\big{(}\mathcal{N}_{\Phi},\langle\cdot,\cdot\rangle_{\mathcal{N}_{\Phi}}\big{)}$. In this context, we will use the function $\Phi$ and its associated kernel $K$ synonymously. A popular class of strictly positive RBFs are the Matérn kernels or Sobolev splines $\Phi_{\theta}\colon\mathbb{R}^{d}\to\mathbb{R}$, dependent on the hyper-parameter $\theta>d/2$, that steers the smoothness of the kernel. These kernels are defined via the radial function
(1) $\phi_{\theta}(r)=\frac{2^{1-\theta+d/2}}{\Gamma(\theta-d/2)}r^{\theta-d/2}K_{\theta-d/2}(r),\quad r\geq 0,$
@@
where $\Gamma$ is the Riemann gamma function and $K_{\nu}$ is the modified Bessel function of the second kind of order $\nu$, see [19] for example. It is well-known that the Fourier transform of this radial function satisfies
(2) $\mathcal{F}{\Phi}(\boldsymbol{\omega})=(1+\|\boldsymbol{\omega}\|_{2}^{2})^{-\theta},\quad\boldsymbol{\omega}\in\mathbb{R}^{d},$
@@
and that, in turn, the native space of the corresponding Matérn kernel is norm-equivalent to the Sobolev space $H^{\theta}(\mathbb{R}^{d})$, see [28, Corollary 10.13] for example.
@@
In this instationary setting, meaning that the length-scale of the kernel is not coupled to the fill-distance, it is known that the interpolant $s\in\mathcal{S}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X\}\subset\mathcal{N}_{\Phi}$ converges towards the function $f$ as $h_{X,\Omega}\to 0$. However, this improved approximation comes with higher and higher numerical cost for the iterative solution due to the ill-conditioning of the associated generalized Vandermonde matrix. To alleviate this cost, we can rescale the kernel, i.e., we introduce the parameter $\delta>0$ and define
@@
In this stationary setting, we couple $\delta$ proportional to $h_{X,\Omega}$, i.e., there is a coupling parameter $\eta>0$ such that
(4) $\delta=\eta h_{X,\Omega}.$
In this case, we can control the numerical costs to compute the interpolant but cannot expect convergence. This is known as the trade-off principle.
@@
where we refer to the subscript of $X_{\ell}$ as level with cardinalities $N_{\ell}\mathrel{\mathrel{\mathop{:}}=}\#X_{\ell}$ and fill-distances $h_{\ell}\mathrel{\mathrel{\mathop{:}}=}h_{X_{\ell},\Omega}$. Although in many applications the sets will be nested, for the multiscale method itself, this is not necessary. We assume that the cardinalities of the sets $X_{\ell}$ are increasing with the level, i.e., $N_{\ell}\leq N_{\ell+1}$ or, in other words, $h_{\ell}\geq h_{\ell+1}$. We quantify this relation in Theorem 2.1. In what follows, we consider the stationary setting at each level $\ell$. This yields level-dependent parameters $\delta_{\ell}>0$ and level-dependently rescaled kernels $K_{\ell}$ given by
(6) $K_{\ell}(\boldsymbol{x},\boldsymbol{y})\mathrel{\mathrel{\mathop{:}}=}\Phi_{\ell}(\boldsymbol{x}-\boldsymbol{y})=\delta_{\ell}^{-d}\Phi\left(\frac{\boldsymbol{x}-\boldsymbol{y}}{\delta_{\ell}}\right),\quad\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d},\ \ell=1,\ldots,L.$
@@
called the local approximation space.
Now, given the multiscale sequence $X_{1},\ldots,X_{L}$, we seek an approximation in terms of successive residual corrections. To this end, we introduce functions $w_{\ell}$ of the form
(8) $w_{\ell}=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i})\in\mathcal{S}_{\ell}.$
The coefficents $c_{\ell,i}$ are determined from decomposing the approximation problem by a telescoping sum according to
@@
This means that, at level $\ell$, we compute the interpolant to the residual at level $\ell-1$, i.e.,
(9) $f-(w_{1}+\cdots+w_{\ell-1})$
in the space $\mathcal{S}_{\ell}$. From this, we obtain the multiscale decomposition $s_{L}\mathrel{\mathrel{\mathop{:}}=}(w_{1}+\cdots+w_{L})\in\cup_{\ell=1}^{L}\mathcal{S}_{\ell}$ which matches $f$ at each of the data sites in $X$, i.e., $s_{L}|_{X}=f|_{X}$.
In practice, the multiscale approximation process begins with a small set of points, where the approximation problem (8) is solved. Next, the error function (9) is computed, and finally, the latter is approximated at a finer scale. The coefficients of the residual correction at level $\ell$ can be computed by solving
$\boldsymbol{K}_{\ell,\ell}\boldsymbol{c}_{\ell}=\boldsymbol{f}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}_{\ell,\ell^{\prime}}\boldsymbol{c}_{\ell^{\prime}}$
with the generalized Vandermonde matrices $\boldsymbol{K}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\big{[}\Phi_{\ell^{\prime}}\big{(}{\boldsymbol{x}}_{i}^{(\ell)}-\boldsymbol{x}_{j}^{(\ell^{\prime})}\big{)}\big{]}_{i,j}\in\mathbb{R}^{N_{\ell}\times N_{\ell^{\prime}}}$ and the vectors $\boldsymbol{f}_{\ell}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{f}|_{X_{\ell}}\in\mathbb{R}^{N_{\ell}},\ \boldsymbol{c}_{\ell}\mathrel{\mathrel{\mathop{:}}=}[{c}_{\ell,i}]_{i}\in\mathbb{R}^{N_{\ell}}.$ Consequently, the coefficient vectors ${\boldsymbol{c}}_{1},\ldots,{\boldsymbol{c}}_{L}$ are obtained by solving the block lower triangular system
(10) $\mathcal{K}{\boldsymbol{c}}\mathrel{\mathrel{\mathop{:}}=}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]\mathrel{=\mathrel{\mathop{:}}}{\boldsymbol{f}}.$
The resulting block matrix $\mathcal{K}$ is in general a densely-populated lower triangular matrix as a result of the choice of globally supported RBFs for the approximation at each level. There holds the following convergence result, see, e.g., [29, 30].
@@
for $\ell=1,2,\dots$ with fixed $\mu\in(0,1)$, $c\in(0,1]$ and $h_{1}$ sufficiently small. Further, let $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel for $H^{\theta}(\mathbb{R}^{d})$, i.e., its Fourier transform satisfies $\mathcal{F}{\Phi}({\boldsymbol{\omega}})\sim(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-\theta},\ {\boldsymbol{\omega}}\in\mathbb{R}^{d}$ and let $\Phi_{\ell}$ be the rescaled kernel as in (3) with scale factor $\delta_{\ell}=\eta h_{\ell}$, $\eta>1$. Let $f\in H^{\theta}(\Omega)$. Then there exist constants $C$ and $C_{1}>0$ such that
$\|f-s_{L}\|_{L_{2}(\Omega)}\leq C(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{L}\|f\|_{H^{\theta}(\Omega)}.$